# Local Loss Optimization in the Infinite Width: Stable Parameterization of Predictive Coding Networks and Target Propagation

### v1

@@ 5.1 $\mu$P of TP
Figure 6: In Target Prop, using $\mu$P ensures that $\Delta h_{l}$ remains consistent across widths. This figure shows the RMS norm of $\Delta h_{l}$ during training. For SP, $\Delta h_{l}$ in the input layer diminishes as the width increases, while $\Delta h_{l}$ in the output layer diverges with increasing width. Consequently, the training dynamics become unstable. In contrast, with $\mu$P, $\Delta h_{l}$ remains consistent across different widths in both the input and output layers.
As overviewed in Section 3.1.2, TP reduces to GNT in the highly restrictive case of invertible networks. However, TP is not equivalent to GNT or BP in general cases [Meulemans et al., 2020, Ernoult et al., 2022]. While TP involves two networks trained using different manners, and one may feel it challenging to obtain a stable parameterization for learning, we demonstrate that, under the assumption that the feedback network uses a linear activation function $\psi$, we can systematically derive $\mu$P for both TP and DTP.
@@
Consider a linear feedback network. The forward network is allowed to have nonlinear activation functions. After the first training phase of $Q_{l}$, take the first one-step update of $W$. Then, we obtain the $\mu$P as follows:
@@
The derivation is presented in Section A.4.1. Note that the linear feedback network has trained weights in a pseudo-inverse form, that is, $Q_{l}^{*}=h_{l-1}(h_{l}^{\top}h_{l}+\mu I)^{-1}h_{l}^{\top}$. Stable parametrization can also be discussed for the training of the feedback network. For further details, see Section A.5.
Figure 6 tracks $\Delta h_{l}$ during training. In $\mu$P, $\Delta h_{l}$ remains consistent across different widths, whereas in SP, $\Delta h_{l}$ either diverges or diminishes as the width changes. As demonstrated in Figure 7, using the $\mu$P for TP results in the $\mu$Transfer appropriately across widths.
@@
Figure 7: (Left) $\mu$P can transfer the learning rate across widths in TP. We trained a 3-layer MLP on FashionMNIST. (Right) TP does not have kernel regime. We measured $\omega_{L}=\log_{M}(\|W_{L,0}\Delta h_{L-1,T}\|_{\text{RMS}}/{\|W_{L,0}\|_{\text{RMS}}\|\Delta h_{L-1,T}\|_{\text{RMS}}})$ across different parametrizations following Everett et al. [2024]. In the infinite-width limit, $\omega_{L}$ converges to $\alpha$. Therefore, in TP, where $\omega_{L}$ remains fixed at 1/2 even as the width increases, the kernel regime disappears.
@@ 5.2 Disappearance of the Kernel Regime
It is notable that $\mu$P in the previous work of the gradient methods requires $b_{L}=1$; in TP, $\mu$P requires $b_{L}=1/2$. For the usual gradient methods, a stable parametrization with $b_{L}=1/2$ leads to the kernel regime. This raises the question: does a kernel regime exist in TP? Interestingly, in TP, the kernel regime disappears (see Corollary A.8 for the details).
@@
It is known that when the inner product $W_{L}\Delta h_{L-1}$ follows the Law of Large Numbers (LLN), $\alpha=1$, and when it follows the Central Limit Theorem (CLT), $\alpha=1/2$ [Everett et al., 2024]. Additionally, to prevent the output of the last layer from exploding, it is necessary that $h_{L}=O(1)$, that is, $a_{L}+b_{L}\geq 1/2$. Consequently, $r\leq\alpha-1/2$. In BP, the dependence between $W_{L}$ and $\Delta h_{L-1}$ results in $\alpha=1$ by the LLN. We have $r\leq 1/2$, allowing for the kernel regime. In contrast, in TP, updating the feedforward network weights does not induce a dependence between $W_{L}$ and $\Delta h_{L-1}$, leading to $\alpha=1/2$. This is because the gradient propagated from the last layer is computed based on the feedback weight $Q^{*}_{L}=h_{L-1}\left(h_{L}^{\top}h_{L}+\mu I\right)^{-1}h_{L}^{\top}$. This corresponds to a regression of the last layer by using the downstream layer as the basis. Consequently, $r\leq 0$ and the kernel regime cannot be achieved in TP.
Note that $\alpha=1/2$ in TP is also empirically confirmed in Figure 7. TP seems to be the first example in the infinite-width limit where $b_{L}=1/2$ induces feature learning.
@@ 6 Conclusion
The derivation of $\mu$P assumes a one-step gradient and linear networks, although this prerequisite is not unique to our work [Yang, 2020, Yang et al., 2024]. Ensuring the existence of feature learning dynamics for more general steps in the infinite width limit would require the development of a tensor program. However, handling the dependencies between variables that differ from standard BP, such as those arising from the inference phase and feedback pass, is non-trivial and presents an interesting direction for future research. Additionally, it would also be valuable to explore the learning dynamics of local learning and its convergence properties by extending the infinite width theory or further analyzing of linear networks. We believe that understanding the universal behavior of large-scale limits will provide a foundation for the development of more effective algorithms.
@@
Appendix A Derivation of Stable Parameterization
A.1 Conditions
@@
We follow the derivation based on the infinitesimal one-step update from random initialization [Yang and Hu, 2021, Ishikawa and Karakida, 2024], which involves taking the limit of a sufficiently small coefficient of the learning rate $\eta^{\prime}$. This formulation clarifies the proof and enables the systematic derivation of $\mu$P across various problems. In the infinitesimal formulation, Conditions 3.1 and 3.2 are expressed as follows:
Condition A.1 ($W_{l}$ updated maximally).
@@
Condition A.2 ($W_{L}$ initialized maximally).
@@
In addition, the stability of learning is defined as follows (see Definition H.4 in Yang and Hu [2021] for more detail):
Definition A.3 (Stability of learning).
$\Delta h_{l,t}=O(1).$ (S.3)
As described in the previous work [Yang and Hu, 2021, Ishikawa and Karakida, 2024], Condition 3.1 (or A.1) is required from the expansion of Eq. (9) by the parameter update, yielding
@@
and, in particular, feature learning is characterized by $r_{l}=0$. Condition 3.2 (or S.2) is required to eliminate a uninteresting case in which the hidden layer provides no contribution to the network output. Both NTK and feature learning regimes are characterized by this condition.
Assumption A.4.
$u_{l,0},h_{l,0}=\Theta(1)$   ($l<L$),  $f_{0}=u_{L,0}=O(1)$.
As shown in Theorem H.6 of Yang and Hu [2021], this assumption immediately leads to
A.2 Predictive Coding with single-shot SI
Theorem A.5 (Stable parameterization for PC).
Set inference step sizes $\gamma_{l<L}=\Theta(1)$ and $\gamma_{L}=\gamma^{\prime}/M^{\bar{\gamma}_{L}}$ with a positive constant $\gamma^{\prime}$. Suppose F-initialization and single-shot sequential inference, and consider one-step update of parameters after the inference. For infinitesimal step sizes $\gamma_{L}^{\prime}$ and $\eta^{\prime}$, PC admits the $\mu$P for feature learning at

### v2

@@ 5.1 $\mu$P of TP
As overviewed in Section 3.1.2, TP reduces to GNT in the highly restrictive case of invertible networks. However, TP is not equivalent to GNT or BP in general cases (Meulemans et al., 2020; Ernoult et al., 2022). While TP involves two networks trained using different manners, and one may feel it challenging to obtain a stable parameterization for learning, we demonstrate that, under the assumption that the feedback network uses a linear activation function $\psi$, we can systematically derive $\mu$P for both TP and DTP.
@@
Consider a linear feedback network. The forward network is allowed to have nonlinear activation functions. After the first training phase of $Q_{l}$, take the first one-step update of $W$. Then, we obtain $\mu$P as follows:
@@
The derivation is presented in Section C.1. Note that the linear feedback network has trained weights in a pseudo-inverse form, that is, $Q_{l}^{*}=h_{l-1}(h_{l}^{\top}h_{l}+\mu I)^{-1}h_{l}^{\top}$. Stable parameterization can also be discussed for the training of the feedback network. For further details, see Section C.2.
As demonstrated in Figure 6, using the $\mu$P for TP results in the $\mu$Transfer appropriately across widths. Furthremore, Figure S.14 in Appendix tracks $\Delta h_{l}$ during training. In $\mu$P, $\Delta h_{l}$ remains consistent across different widths, whereas in SP, $\Delta h_{l}$ either diverges or diminishes as the width changes.
@@
Figure 6: (Left) $\mu$P can transfer the learning rate across widths in TP. (Right) TP does not have kernel regime. We measured $\omega_{L}=\log_{M}(\|W_{L,0}\Delta h_{L-1,T}\|_{\text{RMS}}/{\|W_{L,0}\|_{\text{RMS}}\|\Delta h_{L-1,T}\|_{\text{RMS}}})$ across different parameterizations following Everett et al. (2024). In the infinite-width limit, $\omega_{L}$ converges to $\alpha$. Therefore, in TP, where $\omega_{L}$ remains fixed at 1/2 even as the width increases, the kernel regime disappears.
@@ 5.2 Disappearance of the Kernel Regime
It is notable that $\mu$P in the previous work on the gradient methods requires $b_{L}=1$; in TP, $\mu$P requires $b_{L}=1/2$. For the usual gradient methods, a stable parameterization with $b_{L}=1/2$ leads to the kernel regime. This raises the question: does a kernel regime exist in TP? Interestingly, in TP, the kernel regime disappears (see Corollary C.1 for the details).
@@
When the inner product $W_{L}\Delta h_{L-1}$ follows the Law of Large Numbers (LLN), $\alpha=1$, and when it follows the Central Limit Theorem, $\alpha=1/2$ (Everett et al., 2024). Additionally, to prevent the output of the last layer from exploding, it is necessary that $h_{L}=O(1)$, that is, $a_{L}+b_{L}\geq 1/2$. Consequently, $r\leq\alpha-1/2$. In BP, the dependence between $W_{L}$ and $\Delta h_{L-1}$ results in $\alpha=1$ by the LLN. We have $r\leq 1/2$, allowing for the kernel regime. In contrast, in TP, updating the feedforward network weights does not induce a dependence between $W_{L}$ and $\Delta h_{L-1}$, leading to $\alpha=1/2$. This is because the gradient is computed based on the feedback weight $Q^{*}_{L}=h_{L-1}(h_{L}^{\top}h_{L}+\mu I)^{-1}h_{L}^{\top}$, rather than $W_{L}$. Consequently, $r\leq 0$ and the kernel regime cannot be achieved in TP.
Figure 6 empirically confirms $\alpha=1/2$ in TP. TP seems to be the first example in the infinite-width limit where $b_{L}=1/2$ induces feature learning.
@@ 6 Conclusion
The derivation of $\mu$P assumes a one-step gradient and linear networks, although this prerequisite is not unique to our work (Yang & Hu, 2021; Yang et al., 2024). Ensuring the existence of feature learning dynamics for more general steps in the infinite width limit would require the development of a tensor program. However, handling the dependencies between variables that differ from standard BP, such as those arising from the inference phase and feedback pass, is non-trivial and presents an interesting direction for future research. Additionally, it would also be valuable to explore the learning dynamics of local learning and its convergence properties by extending the infinite width theory or further analyzing linear networks. We believe that understanding the universal behavior of large-scale limits will provide a foundation for the development of more effective algorithms.
Acknowledgments
This work is supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAEKENHI (Grant Nos. 22H05116, 23K16965). S.I. and R.Y. also acknowledge support from JST CREST Grant Number JPMJCR2112.
@@ Acknowledgments
Appendix A Extended Background
A.1 Extended related work
Table S.1 : List of Abbreviations
Abbreviation Full Name Reference
BP BackPropagation -
SGD Stochastic Gradient Descent -
GNT Gauss Newton Target Proposition 3.3
$\mu$P Maximal Update Parameterization Definition A.5
NTK Neural Tangent Kernel -
HP HyperParameter -
PC Predictive Coding Section 3.1.1
F-ini Forward initialization Technique i
FPA Fixed prediction assumption Technique ii
TP Target Propagation Section 3.1.2
DTP Difference Target Propagation Eq S.13
A.1.1 $\mu$P
While $\mu$P was introduced as a parameterization to induce a feature-learning regime in the infinite-width limit for theoretical interest (Yang & Hu, 2021), one practical advantage highlighted by Yang et al. (2021) is its ability to transfer the learning rate across different widths, a phenomenon they experimentally validated. This phenomenon, known as $\mu$Transfer, has also been examined from a theoretical perspective (Noci et al., 2024). Another notable advantage of $\mu$P is the improvement in the scaling law exponent, which has been investigated both experimentally (Qiu et al., 2024) and theoretically (Bordelon et al., 2024).
It is noteworthy that $\mu$P depends on the learning algorithm used and thus should be derived for each specific method. The $\mu$P for Adam was introduced in Yang et al. (2021), with its theoretical justification provided in Yang & Littwin (2023). For the second-order optimization, including the Gauss-Newton algorithm, K-FAC, and Shampoo, the $\mu$P was derived in Ishikawa & Karakida (2024). These works emphasize the importance of adjusting not only the learning rate but also the damping term in second-order optimization using $\mu$P. Additionally, $\mu$P was derived for Adafactor in Everett et al. (2024) and empirically demonstrated that the scaling of the $\epsilon$ term in Adam is also crucial in $\mu$P.
A.1.2 Predictive Coding
Recent progress in deep learning has largely been achieved by the success of backpropagation (Rumelhart et al., 1986; LeCun et al., 1998; 2015). This success has increased the interest in exploring whether deep networks can also be trained using training algorithms other than backpropagation. This includes exploration into biologically plausible training methods and the benchmarking of local learning rules at modern-scale networks on deep-learning benchmark datasets; equilibrium propagation (Scellier & Bengio, 2017; Laborieux et al., 2021), target propagation (Bengio, 2014; Lee et al., 2015), predictive coding (Whittington & Bogacz, 2017; Song et al., 2020; Salvatori et al., 2023), and forward-forward algorithms (Ren et al., 2023).
Computation of predictive coding for supervised learning: Predictive coding was originally introduced as an algorithm for solving inverse problems where the goal is to find the parameters $W$ that maximize the marginal likelihood $p(v_{L};W)$ where $v_{L}$ denotes a variable representing the network output and $v_{L}=x$ for the input data $x$ (Rao & Ballard, 1999). In this inverse problem, we consider latent variables $v_{l}$ (also referred to as causes) and a generative function $p(v_{l}|v_{l-1};W_{l})$. If we assume a hierarchical Gaussian generative model, the marginal probability of the causes is as follows:
$p\left(v_{0},\ldots,v_{L};W\right)$ $=p(v_{0})\prod_{l=1}^{L}p(v_{l}|v_{l-1};W_{l})$ (S.1)
$=\mathcal{N}\left(v_{0},\gamma_{0}^{-1}I\right)\prod_{l=1}^{L}\mathcal{N}\left(W_{l}\phi(v_{l-1}),\gamma_{l}^{-1}I\right).$ (S.2)
Friston (2003) framed this inverse problem as an EM algorithm aimed at minimizing the following variational free energy:
$\mathcal{F}=\operatorname{KL}(q(v_{0},\ldots,v_{L})\|p(v_{0},\ldots,v_{L};W)),$ (S.3)
where $q(v_{0},\ldots,v_{L})$ is a tractable posterior probability distribution for the EM algorithm. In particular, the E-step, referred to as inference, minimizes variational free energy by causes $v_{l}$, while the M-step, referred to as learning, minimizes by parameters $W_{l}$. We usually apply a mean-field approximation or a Laplace approximation to the tractable probability distribution (Friston, 2005; Salvatori et al., 2024a; b). Under these formulations, we can derive the variational free energy for PC.
$\mathcal{F}=\sum_{l=1}^{L}\gamma_{l}\frac{1}{2}\left\|v_{l}-W_{l}\phi\left(v_{l-1}\right)\right\|^{2}.$ (S.4)
While predictive coding networks were originally discussed primarily in the context of generative models for unsupervised learning, Whittington & Bogacz (2017) reformulated PC networks for supervised learning and highlighted their potential for use in the context of deep learning. Specifically, if we fix $v_{0}=x$ and $v_{L}=y$ for data, this corresponds to supervised learning using the mean squared loss.
Heuristic techniques in PC: After Whittington & Bogacz (2017), there has been an increasing amount of studies evaluating the performance of PC networks as local learning on deep-learning benchmark datasets (Salvatori et al., 2023; Pinchetti et al., 2024). They revealed that the original implementation of PC networks is insufficient for achieving stable training performance, and heuristic modifications have played an important role. For instance, Fixed prediction assumption (FPA) has been introduced to achieve higher performance by approximating the gradient computation of PC networks closer to that of BP (Millidge et al., 2022b; Rosenbaum, 2022). Under FPA, $\phi^{\prime}\left(v_{l,s}\right)$ is replaced with $\phi^{\prime}\left(v_{l,0}\right)$, resulting in an inference phase given by
$v_{l,s+1}$ $=v_{l,s}-\gamma_{l}e_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,0}\right)\circ W_{l+1}^{\top}e_{l+1,s},\quad(l<L)$ (S.5)
and a learning phase is given by
$W_{l,t+1}$ $=W_{l,t}+\eta_{l}\gamma_{l}e_{l,s}\phi(v_{l-1,0})^{\top}.$ (S.6)
Additionally, FPA is typically used with Forward initialization (F-ini). In Forward initialization, the state $v_{l,0}$ is initialized with the forward value $u_{l,0}$. While F-ini has been implicitly utilized in most studies on PC (Whittington & Bogacz, 2017; Song et al., 2020; Rosenbaum, 2022), its role was explicitly highlighted in Alonso et al. (2022), where the authors compared the convergence with and without F-ini.
Rosenbaum (2022) pointed out that when both F-ini and FPA are assumed, PC networks are entirely reduced to BP and that if the algorithms are fully equivalent to BP, the advantages of biological plausibility and local updates are lost. In this study, we aim to identify a parameterization that enables stable local learning while maintaining distinctions from BP. By leveraging several recently developed heuristics, we clarify the desirable scales for stable and efficient local learning.
Nudged PC: The design of loss functions in PC networks has also been a focus of algorithmic improvements. In most of the ML research, classification tasks generally use cross-entropy loss rather than mean squared loss. Accordingly, PC networks sometimes use cross-entropy loss as well (Pinchetti et al., 2024). The free energy for a general loss function is given by
$\mathcal{F}=\mathcal{L}(y,W_{L}\phi(v_{L-1}))+\sum_{l=1}^{L-1}\gamma_{l}\frac{1}{2}\left\|v_{l}-W_{l}\phi\left(v_{l-1}\right)\right\|^{2}.$ (S.7)
and its update rule for the inference phase is given by
$v_{l,s+1}$ $=v_{l,s}-\gamma_{l}e_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,s}\right)\circ W_{l+1}^{\top}e_{l+1,s},\quad(l<L)$ (S.8)
$v_{L,s+1}$ $=v_{L,s}-\gamma_{L}\frac{\partial\mathcal{L}}{\partial v_{L}}.$ (S.9)
Furthermore, there are formulations of PC networks that incorporate the nudge term introduced in equilibrium propagation (Scellier & Bengio, 2017). PC networks with a nudge term update the state $v_{l}$ and weights $W_{l}$ to minimize the following free energy function (Alonso et al., 2022; Millidge et al., 2023; Pinchetti et al., 2024):
Here, $\beta$ is a nudge coefficient parameter that covers some variants of the PC algorithms in the previous work.
There are several possible orders for computing this inference (Alonso et al., 2024) as illustrated in Algorithm 1. Specifically, $e_{l}$ can be calculated sequentially from the output layer to the input layer, or it can be updated synchronously across all layers simultaneously from the output to the input. While the main text focuses on Sequential Inference (SI), where computations proceed layer-by-layer from the output to the input, Predictive Coding with synchronous inference is also a valid approach worth considering.
It is worth noting that layer-wise learning methods with auxiliary variables, such as the method of auxiliary coordinates (MAC) (Carreira-Perpiñán & Wang, 2014), alternating direction method of multipliers (ADMM) (Taylor et al., 2016), and online alternating minimization (AM) (Choromanska et al., 2019), utilize local losses identical or similar to those in PC. Thus, we expect our $\mu$P framework to be equally effective when applied to these methods.
Algorithm 1 PC Algorithm (Simultaneous or Sequential inference)
1:for $s=1$ to $n$ do
3:     for $l=L-1$ to $1$ do
4:         $e_{l,s}\leftarrow u_{l,s}-W_{l}\phi_{l-1}(u_{l-1,s})$
7:     end for
8:end for
A.1.3 Target Propagation
Target propagation offers a learning rule that is more biologically plausible and easier for the brain to implement compared to BP (Bengio, 2014). Specifically, it addresses the following two issues inherent to BP (Meulemans et al., 2020).
  1. 1.
    Signed error transmission problem: BP propagates the error gradient to the lower layers, whereas the brain propagates target values for the neurons (Lillicrap et al., 2020).
  2. 2.
    Weight transport problem: BP requires exact weight symmetry between the forward and backward paths. However, the brain cannot transport weights (Grossberg, 1987; Akrout et al., 2019).
Target propagation aims to address the two issues by:
  1. 1.
    Propagating the target value $\hat{h}_{L}=h_{L}-\hat{\eta}\nabla_{h_{L}}\mathcal{L}$ instead of the error gradient $\nabla_{h_{L}}\mathcal{L}$ for signed error transmission problem.
  2. 2.
    Utilizing a feedback network $g_{l}$ distinct from the feedforward network $f_{l}$ to propagate the target value $\hat{h}_{L}$ for weight transport problem.
The feedback network $g_{l}$ has the weights $Q_{l}$ distinct from those in the feedforward network.
$\hat{h}_{l}=g_{l}(\hat{h}_{l+1}),\quad g_{l}(x)=\psi(Q_{l}x)\quad(l=1,...,L-1).$ (S.11)
Here, $\psi$ denotes the activation function. While it is often the same as the activation function used in the feedforward network, it is also possible to consider a different activation function. We set $\psi=\phi$ in experiments and assume $\psi$ as the identity function only in Theorem 5.1. The feedback network is trained to minimize the following reconstruction loss:
$\mathcal{L}_{\textbf{rec}}(Q_{l})=\left\|g_{l}\left(f_{l}(h_{l-1}+\epsilon)\right)-h_{l-1}+\epsilon\right\|^{2},$ (S.12)
where $\epsilon$ is a small Gaussian noise to improve the robustness of the feedback network. Target propagation attempts to approximate the inverse function of the feedforward network by learning the feedback network through the optimization of this reconstruction loss.
Difference target propagation (DTP) is an improved method of TP, which adjusts the propagation in the feedback network as follows (Lee et al., 2015).
In TP, the accumulation of the reconstruction error $g_{i}(f_{i}(h_{i}))-h_{i}$ during propagation poses an obstacle to optimization. In DTP, subtracting $\left(g_{i}\left(h_{i+1}\right)-h_{i}\right)$ mitigates the accumulation of the reconstruction error and improves the progress of learning.
As a side note, Meulemans et al. (2020) and Bengio (2020) pointed out that TP can be related to the Gauss-Newton method for invertible networks. Additionally, Meulemans et al. (2020) proposed Direct Difference Target Propagation so as to establish this correspondence even in non-invertible networks under some infinitesimal conditions. Ernoult et al. (2022) reported that one can stabilize TP by introducing the additional Local Difference Reconstruction Loss which makes the gradient align more closely with Backpropagation rather than Gauss-Newton Targets. In our work, we aim to clarify the fundamental properties of TP and DTP from the perspective of parameterization and do not consider such additional conditions or loss functions.
A.2 Definitions for stable parameterization
Assumption A.1.
$u_{l,0},h_{l,0}=\Theta(1)$   ($l<L$),  $f_{0}=u_{L,0}=O(1)$.
As shown in Theorem H.6 of Yang & Hu (2021), this assumption immediately leads to
In addition, the stability of learning is defined as follows (see Definition H.4 in Yang & Hu (2021) for more detail):
Definition A.2 (Stability of learning).
We say an abc-parameterization is stable if, for $l<L$ and for any fixed $t\geq 1$,
$\Delta h_{l,t}=O(1),\ \Delta f_{t}=O(1),$ (S.15)
under Assumption A.1.
Condition S.15 ensures avoiding exploding dynamics with respect to the width, i.e., $\Delta h_{l,t}=O(1/M^{k})$ with $k<0$.
@@ A.2 Definitions for stable parameterization
We follow the derivation based on the infinitesimal one-step update from random initialization (Yang & Hu, 2021; Ishikawa & Karakida, 2024), which involves taking the limit of a sufficiently small coefficient of the learning rate $\eta^{\prime}$. This formulation clarifies the proof and enables the systematic derivation of $\mu$P across various problems. In the infinitesimal formulation, Conditions 3.1 and 3.2 are expressed as follows:
Condition A.3 ($W_{l}$ updated maximally).
@@
Condition A.4 ($W_{L}$ initialized maximally).
@@
As described in the previous work (Yang & Hu, 2021; Ishikawa & Karakida, 2024), Condition 3.1 (or A.3) naturally appears from the expansion of Eq. (10) by the parameter update, yielding
@@
and, in particular, feature learning is characterized by $r_{l}=0$. Condition 3.2 (or S.17) is required to eliminate an uninteresting case in which the hidden layer provides no contribution to the network output. Both NTK and feature learning regimes are characterized by this condition.
As is shown in Yang & Hu (2021), $\mu$P is the unique stable parameterization satisfying Condition A.3 for $l\leq L$ and Condition S.17 for $W_{L}$. Thus, we can admit this characterization as a definition of $\mu$P.
Definition A.5 ($\mu$P).
$\mu$P is the stable abc parameterization satisfying Condition A.3 for $l\leq L$ and Condition S.17 for $W_{L}$.
Note that Condition A.3 is required not only for hidden layers but also for the last layer. In the previous work, this eliminates a trivial case of learning, i.e., $\Delta h_{L,t}=O(1/M^{k})$ with $k>0$, where the effect of learning vanishes.
Appendix B $\mu$P of Predictive Coding
B.1 Derivation for Predictive Coding with single-shot SI
Theorem B.1 (Stable parameterization for PC).
Set inference step sizes $\gamma_{l<L}=\Theta(1)$ and $\gamma_{L}=\gamma^{\prime}/M^{\bar{\gamma}_{L}}$ with a positive constant $\gamma^{\prime}$. Suppose F-ini and single-shot sequential inference, and consider a one-step update of parameters after the inference. For infinitesimal step sizes $\gamma_{L}^{\prime}$ and $\eta^{\prime}$, PC admits the $\mu$P for feature learning at